OpenStudy (anonymous):

How many solutions does the equation 6z – 3z – 7 = –2 + 3 have? Two None Infinitely many One

OpenStudy (anonymous):

is this the cosketch gal?

OpenStudy (anonymous):

i have no clue what you are talking about

OpenStudy (anonymous):

ooops srry nvrmnd but yeah i can help you!

OpenStudy (anonymous):

do you need help on this particular problem?

OpenStudy (anonymous):

yes and more

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

are you there?

OpenStudy (anonymous):

so remember to combine like terms. which combination of terms would you combine first?

OpenStudy (anonymous):

i think it was the 6 and the 3

OpenStudy (anonymous):

ok first of all start with saying 6z and 3z so you dont confuse yourself later on :) but yes do combine those

OpenStudy (anonymous):

i am so confused here!

OpenStudy (anonymous):

yes you were right combine the 6z with the 3z.

OpenStudy (anonymous):

you are not helping me like you were earlier whats wrong

OpenStudy (anonymous):

Try this, www.wolframalpha.com It's a computational knowledge engine.(Basically a very big calculator) You can plug the equation into the engine and you can see the solution. I would suggest getting the app, it's cheap and it can show the step-by-step solution.

OpenStudy (anonymous):

Can you try doing the problem? Like do you know where to start? I remember doing a couple of problems with u and i want to see if u can apply the information i taught u to this problem here :). Sorry i get distracted easily :(

OpenStudy (anonymous):

hey can you tell me if two is right!

OpenStudy (anonymous):

Is the problem written correctly ? Is there only 1 variable in this problem ?

OpenStudy (anonymous):

let me check

OpenStudy (anonymous):

yup!

OpenStudy (anonymous):

6z - 3z - 7 = -2 + 3 ---- combine like terms 3z - 7 = 1 do you understand so far ?

OpenStudy (anonymous):

@agl202309 please try this problem on your own. but if not you need to combine . 6z - 3z

OpenStudy (anonymous):

yes(sorry my computer is very slow)

OpenStudy (anonymous):

add seven on both sides of the equation so that seven cancels out on the left side of the equation so now it would look like this 3z = 1 + 7

OpenStudy (anonymous):

now combine the terms on the right side

OpenStudy (anonymous):

the terms are 7 and 1 so combine them and you get what?

OpenStudy (anonymous):

so it is 8

OpenStudy (anonymous):

no.. we now have 3z = 8 now divide both sides by 3

OpenStudy (anonymous):

it is indeed 8 but it is 3z = 8 now you need to isolate the z which means dividing it on both sides of the equation so now it looks like z = 8/3

OpenStudy (anonymous):

correct....so if z = 8/3.....that is 1 solution

OpenStudy (anonymous):

becuase there are no infinite/none amount of answer s for this one problem.

OpenStudy (anonymous):

ok one solution! thanks for helping.

OpenStudy (anonymous):

YW :)

OpenStudy (anonymous):

np :)

OpenStudy (anonymous):

so what are the other problems?